OpenStudy (gabylovesyou):

Rationalize the denominator of square root of negative 49 over open parentheses 7 minus 2 i close parentheses minus open parentheses 4 plus 9 i.

OpenStudy (gabylovesyou):

\[\frac{ \sqrt{49} }{ (7 - 2i) - (4 + 9i) }\]

OpenStudy (gabylovesyou):

@iGreen @sleepyjess

OpenStudy (anonymous):

\[\left(\begin{matrix}\sqrt{49} \\ -28-63i+8i+18i^2\end{matrix}\right)\]

OpenStudy (gabylovesyou):

@pierodog i simplify the equation ? -46 - 55i

OpenStudy (radar):

simplify the numerator and the denominator getting:

OpenStudy (gabylovesyou):

what fraction ?

OpenStudy (radar):

The fraction you submitted........???

OpenStudy (gabylovesyou):

._. idk how to do it ;(

OpenStudy (radar):

This fraction:\[\sqrt{49}\over (7 - 2i) - (4 + 9i)\] simplify the numerator which becomes what i posted in the drawing simplify the denominator as I also shown, ie perform the subtraction.

OpenStudy (radar):

When you get to, the simplified fraction: